# Sim-to-real causal transfer: A metric learning approach to causally-aware interaction representations

### v1

@@ Abstract
Modeling spatial-temporal interactions among neighboring agents is at the heart of multi-agent problems such as motion forecasting and crowd navigation. Despite notable progress, it remains unclear to which extent modern representations can capture the causal relationships behind agent interactions. In this work, we take an in-depth look at the causal awareness of these representations, from computational formalism to real-world practice. First, we cast doubt on the notion of non-causal robustness studied in the recent CausalAgents benchmark (Roelofs et al., 2022). We show that recent representations are already partially resilient to perturbations of non-causal agents, and yet modeling indirect causal effects involving mediator agents remains challenging. To address this challenge, we introduce a metric learning approach that regularizes latent representations with causal annotations. Our controlled experiments show that this approach not only leads to higher degrees of causal awareness but also yields stronger out-of-distribution robustness. To further operationalize it in practice, we propose a sim-to-real causal transfer method via cross-domain multi-task learning. Experiments on pedestrian datasets show that our method can substantially boost generalization, even in the absence of real-world causal annotations. We hope our work provides a new perspective on the challenges and potential pathways towards causally-aware representations of multi-agent interactions. Our code is available at https://github.com/socialcausality.
@@ 1 Introduction
Modeling multi-agent interactions with deep neural networks has made great strides in the past few years (Alahi et al., 2016; Vemula et al., 2018; Gupta et al., 2018; Kosaraju et al., 2019; Salzmann et al., 2020; Mangalam et al., 2021; Gu et al., 2021; Xu et al., 2022; Chen et al., 2023). Yet, existing representations still face tremendous challenges in handling changing environments: they often suffer from substantial accuracy drops under mild environmental changes (Liu et al., 2021; Bagi et al., 2023) and require a large number of examples for adaptation to new contexts (Moon & Seo, 2022; Kothari et al., 2022). These challenges are arguably rooted in the nature of the learning approach that seeks statistical correlations in the training data, regardless of their stability and reusability across distributions (Castri et al., 2022; Bagi et al., 2023). One promising solution is to build more causally-aware representations – latent representations that are capable of capturing the invariant causal dependencies behind agent interactions.
@@
However, discovering causal knowledge from observational data is often exceptionally difficult (Schölkopf et al., 2021; Schölkopf & von Kügelgen, 2022). Most prior works resort to additional information, such as structural knowledge (Chen et al., 2021; Makansi et al., 2021) and domain labels (Liu et al., 2022; Hu et al., 2022; Bagi et al., 2023). While these attempts have been shown effective in certain out-of-distribution scenarios, they still fall short of explicitly accounting for causal relationships between interactive agents. More recently, CausalAgents (Roelofs et al., 2022) made an effort to collect annotations of agent relations (causal or non-causal) in the Waymo dataset (Ettinger et al., 2021), providing a new benchmark focused on the robustness issue under non-causal agent perturbations. Nevertheless, the reason behind the robustness issue remains unclear, as does the potential use of the collected annotations for representation learning.
@@
The goal of this work is to provide an in-depth analysis of the challenges and opportunities of learning causally-aware representations in the multi-agent context. To this end, we first take a critical look at the recent CausalAgents (Roelofs et al., 2022) benchmark. We find that its labeling procedure and evaluation protocol, unfortunately, present subtle yet critical caveats, thereby resulting in a highly biased measure of robustness. To mitigate these issues, we construct a diagnostic dataset through counterfactual simulations and revisit a collection of recent multi-agent forecasting models. Interestingly, we find that most recent models are already partially resilient to perturbations of non-causal agents but still struggle to capture indirect causal effects that involve mediator agents.
@@
To further enhance the causal robustness of the learned representations, we propose a regularization approach that seeks to preserve the causal effect of each individual neighbor in an embedding space. Specifically, we devise two variants that exploit annotations with different levels of granularity: (i) a contrastive-based regularizer using binary annotations of causal/non-causal agents; (ii) a ranking-based regularizer using continuous annotations of causal effects. Through controlled experiments, we show that both regularizers can enhance causal awareness to notable degrees. More crucially, we find that finer-grained annotations are particularly important for generalization out of the training distribution, such as higher agent density or unseen context arrangements.
@@
Finally, we introduce a sim-to-real causal transfer framework, aiming at extending the strengths of causal regularization from simulation to real-world contexts. We achieve this through cross-domain multi-task learning, i.e., jointly train the representation on the causal task in simulation and the forecasting task in the real world. Through experiments on the ETH-UCY dataset (Lerner et al., 2007; Pellegrini et al., 2010) paired with an ORCA simulator (van den Berg et al., 2008), we find that the causal transfer framework enables stronger generalization in challenging settings such as low-data regimes, even in the absence of real-world causal annotations. As one of the first steps towards causal models in the multi-agent, we hope our work brings new light on the challenges and opportunities of learning causally-aware representations in practice.
Figure 1: Illustration of multi-agent interactions. The behavior of the ego agent is causally influenced by some neighbors. We study the challenges and potential for modeling the causal relations between interactive agents.
@@ 2 Related Work
The study of multi-agent interactions has a long history. Early efforts were focused on hand-crafted rules, such as social forces (Helbing & Molnar, 1998; Mehran et al., 2009) and reciprocal collision avoidance (van den Berg et al., 2008; Alahi et al., 2014). Despite remarkable results in sparse scenarios (Luber et al., 2010; Zanlungo et al., 2011; Ferrer et al., 2013), these models often lack social awareness in more densely populated and complex environments (Rudenko et al., 2020). As an alternative, recent years have witnessed a paradigm shift toward learning-based approaches, particularly the use of carefully designed neural networks to learn representations of multi-agent interactions (Kothari et al., 2021). Examples include pooling operators (Alahi et al., 2016; Gupta et al., 2018; Deo & Trivedi, 2018), attention mechanisms (Vemula et al., 2018; Sadeghian et al., 2019; Huang et al., 2019), spatio-temporal graphs (Kosaraju et al., 2019; Salzmann et al., 2020; Li et al., 2020), among others (Rhinehart et al., 2019; Chai et al., 2020; Choi et al., 2021). Nevertheless, the robustness of these models remains a grand challenge (Saadatnejad et al., 2022; Roelofs et al., 2022; Cao et al., 2023). Our work presents a solution to enhance robustness by effectively exploiting causal annotations of varying granularity.
Robust Representations.
@@
The robustness of machine learning models, especially under distribution shifts, has been a long-standing concern for safety-critical applications (Recht et al., 2019). Existing efforts have explored two main avenues to address this challenge. One line of work seeks to identify features that are invariant across distributions. Unfortunately, this approach often relies on strong assumptions about the underlying shifts (Hendrycks* et al., 2020; Liu et al., 2021) or on access to multiple training domains (Arjovsky et al., 2020; Krueger et al., 2021), which may not be practical in real-world settings. Another approach aims to develop models that can efficiently adapt to new distributions by updating only a small number of weight parameters, such as sub-modules (Kothari et al., 2022), certain layers (Lee et al., 2023), or a small subset of neurons (Chen et al., 2022). More recently, there has been a growing interest in exploring the potential of causal learning to address the robustness challenges (van Steenkiste et al., 2019; Dittadi et al., 2020; Montero et al., 2022; Dittadi et al., 2022). To the best of our knowledge, our work makes the first attempt in the multi-agent context, showcasing the benefits of incorporating causal relationships for stronger out-of-distribution generalization in more crowded spaces.
Causal Learning.
@@
Empowering machine learning with causal reasoning has gained a growing interest in recent years (Schölkopf, 2019; Schölkopf & von Kügelgen, 2022). One line of work seeks to discover high-level causal variables from low-level observations, e.g., disentangled (Chen et al., 2016; Higgins et al., 2017; Locatello et al., 2019) or structured latent representations (Locatello et al., 2020; Schölkopf et al., 2021). Unfortunately, existing methods remain largely limited to simple and static datasets (Liu et al., 2023). Another thread of work attempts to draw causal insights into dynamic decision-making (Huang et al., 2022). In particular, recent works have proposed a couple of methods to incorporate causal invariance and structure into the design and training of forecasting models in the multi-agent context (Chen et al., 2021; Makansi et al., 2021; Liu et al., 2022; Hu et al., 2022; Castri et al., 2022; Bagi et al., 2023). However, these efforts have mainly focused on using causal implications to enhance robustness rather than explicitly examining causal relations between interactive agents. Closely related to ours, another recent study introduces a motion forecasting benchmark with annotations of causal relationships (Roelofs et al., 2022). Our work takes a critical look at its labeling as well as evaluation protocols and designs practical methods to enhance causal awareness.
3 Formalism
@@
In this section, we seek to formalize the robustness challenge in multi-agent representation learning through the lens of social causality. We will first revisit the design of the CausalAgents (Roelofs et al., 2022) benchmark, drawing attention to its caveats in labeling and evaluation. Subsequently, we will introduce a diagnostic dataset, aiming to facilitate more rigorous development and evaluation of causally-aware representations through counterfactual simulations.
@@ 3.1 Social Interaction and Social Causality
Social Interaction.
Consider a motion forecasting problem where an ego agent is surrounded by a set of neighboring agents in a scene. Let $s_{t}^{i}=(x_{t}^{i},y_{t}^{i})$ denote the state of agent $i$ at time $t$ and $s_{t}=\{s_{t}^{0},s_{t}^{1},\cdots,s_{t}^{K}\}$ denote the joint state of all the agents in the scene. Without loss of generality, we index the ego agent as 00 and the rest of neighboring agents as $\mathcal{A}=\{1,2,\cdots,K\}$. Given a sequence of history observations $\bm{x}=(s_{1},\cdots,s_{t})$, the task is to predict the future trajectory of the ego agent $\bm{y}=(s_{t+1}^{0},\cdots,s_{T}^{0})$ until time $T$. Modern forecasting models are largely composed of encoder-decoder neural networks, where the encoder $f(\cdot)$ first extracts a compact representation $\bm{z}$ of the input with respect to the ego agent and the decoder $g(\cdot)$ subsequently rolls out a sequence of its future trajectory $\bm{\hat{y}}$:
@@
Social Causality.
Despite remarkable progress on accuracy measures (Alahi et al., 2016; Salzmann et al., 2020), recent neural representations of social interactions still suffer from a significant robustness concern. For instance, recent works have shown that trajectories predicted by existing models often output colliding trajectories (Liu et al., 2021), vulnerable to adversarial perturbations (Saadatnejad et al., 2022; Cao et al., 2022) and deteriorate under distribution shifts of spurious feature such as agent density (Liu et al., 2022).
One particular notion of robustness that has recently gained much attention is related to the causal relationships between interactive agents (Roelofs et al., 2022). Ideally, neural representations of social interactions should capture the influence of neighboring agents, namely how the future trajectory of the ego agent will vary between a pair of scenes: an original scene $\bm{x}_{\varnothing}$ and a perturbed scene $\bm{x}_{R}$ where some neighboring agents $\mathcal{R}$ are removed, i.e., only the ego agent and $\mathcal{A}\setminus\mathcal{R}$ remain. We define the causal effect of the removed agents $\mathcal{R}$ as,
@@
where $\bm{y}_{\varnothing}\equiv\bm{y}$ is the future trajectory in the original scene, $\bm{y}_{\mathcal{R}}$ is the future trajectory in the perturbed scene, and $\|\cdot\|_{2}$ is the average point-wise Euclidean distance between two trajectories.
3.2 Caveats of Social Causality Benchmark
While the mathematical definition of the causal effect is straightforward in the multi-agent context, measuring Eq. 2 in practice can be highly difficult. In general, it is impossible to observe a subset of agents replaying their behaviors in the same environment twice in the real world – an issue known as the impossibility of counterfactuals (Peters et al., 2017). To mitigate the data collection difficulty, a recent benchmark CausalAgents (Roelofs et al., 2022) proposes another simplified labeling strategy: instead of collecting paired counterfactual scenes, it queries human labelers to divide neighboring agents into two categories: causal agents that directly influence the driving behavior of the ego agent from camera viewpoints; and non-causal agents that do not. This labeling approach is accompanied by an evaluation protocol through agent removal, assuming that robust forecasting models should be insensitive to scene-level perturbations that remove non-causal agents. More formally, the CausalAgents benchmark evaluates the robustness of a learned representation through the following measure,
@@ 3.2 Caveats of Social Causality Benchmark
where $\mathcal{R}$ is the set of all (or some) non-causal agents in the scene.
[Uncaptioned image] Figure 2: Illustration of indirect causal effects. The cyclist indirectly influences the decision of the ego agent due to the presence of the bus.
[Uncaptioned image] Figure 3: Joint effect of all non-causal agents in the scene, simulated in ORCA. The ego agent often changes its behavior when $2+$ non-causal neighbors are removed.
Caveats.
In spite of the precious efforts on human annotations, the CausalAgents benchmark (Roelofs et al., 2022) is unfortunately plagued by two subtle but fundamental flaws.
  1. 1.
    Annotation: The labeling rule completely overlooks indirect causal relationships, i.e., a neighbor does not directly influence the behavior of the ego agent but does so indirectly by influencing one or a few other neighbors that pass the influence to the ego agent, as illustrated in Fig. 2. On the one hand, indirect causal effects are practically non-trivial for human labelers to identify due to complex relation chains. On the other hand, they are prevalent in densely populated scenes, not only posing a significant modeling challenge but also playing a crucial role in causal learning, which we will demonstrate in Sec. 5.
  2. 2.
    Evaluation: The evaluation protocol tends to dramatically overestimate the robustness issue. As per Eq. 3, non-causal agents are delineated at the individual level, but robustness assessment occurs at the category level, e.g., removing all non-causal agents rather than a single one. In fact, the joint effect of non-causal agents can escalate quickly with an increasing number of agents, as demonstrated in Fig. 3. Such a discrepancy between annotation and evaluation with respect to non-causal agents can lead to an inflated perception of the robustness issue.

### v2

@@ Abstract
Modeling spatial-temporal interactions among neighboring agents is at the heart of multi-agent problems such as motion forecasting and crowd navigation. Despite notable progress, it remains unclear to which extent modern representations can capture the causal relationships behind agent interactions. In this work, we take an in-depth look at the causal awareness of these representations, from computational formalism to real-world practice. First, we revisit the notion of non-causal robustness studied in the recent CausalAgents benchmark [56]. We show that existing representations are already partially resilient to perturbations of non-causal agents, and yet modeling indirect causal effects involving mediator agents remains challenging. To address this challenge, we introduce a metric learning approach that regularizes latent representations with causal annotations. Our controlled experiments show that this approach not only leads to higher degrees of causal awareness but also yields stronger out-of-distribution robustness. To further operationalize it in practice, we propose a sim-to-real causal transfer method via cross-domain multi-task learning. Experiments on trajectory prediction datasets show that our method can significantly boost generalization, even in the absence of real-world causal annotations, where we acquire higher prediction accuracy by only using 25% of real-world data. We hope our work provides a new perspective on the challenges and pathways toward causally-aware representations of multi-agent interactions. Code is available at https://github.com/vita-epfl/CausalSim2Real.
@@ 1 Introduction
Modeling multi-agent interactions with deep neural networks has made great strides in the past few years [2, 68, 24, 31, 61, 46, 23, 69, 11]. Yet, existing representations still face tremendous challenges in handling changing environments: they often suffer from substantial accuracy drops under mild environmental changes [39, 4] and require a large number of examples for adaptation to new contexts [51, 33]. These challenges are arguably rooted in the nature of the learning approach that seeks statistical correlations in the training data, regardless of their stability and reusability across distributions [7, 4]. One promising approach is to develop causally-aware representations—latent features that capture the invariant causal dependencies in agent interactions.
Figure 1: Illustration of causal relations in multi-agent interactions. The behavior of the ego agent is causally influenced by some neighboring agents. Our research explores the depth of causal understanding within trajectory prediction models, introduces novel causal regularizers to refine this understanding, and details a sim-to-real approach that effectively transfers this causal knowledge to practical real-world scenarios.
@@
However, discovering causal knowledge from observational data is often exceptionally difficult [64, 63]. Most prior works resort to additional information, such as structural knowledge [9, 45] and domain labels [40, 28, 4]. While these attempts have been shown effective in certain out-of-distribution scenarios, they still fall short of explicitly accounting for causal relationships between interactive agents. More recently, CausalAgents [56] made an effort to collect annotations of agent relations (causal or non-causal) in the Waymo dataset [20], providing a new benchmark focused on the robustness issue under non-causal agent perturbations. Nevertheless, the reason behind the robustness issue remains unclear, as does the potential use of the collected annotations for representation learning.
@@
The goal of this work is to provide an in-depth analysis of the challenges and opportunities of learning causally-aware representations in the multi-agent context. To this end, we first take an in-depth look at the recent CausalAgents benchmark [56]. We find that its labeling procedure and evaluation protocol present subtle yet critical caveats, thereby resulting in a highly biased measure of robustness. To mitigate these issues, we construct a diagnostic dataset through counterfactual simulations and revisit a collection of recent multi-agent forecasting models. Interestingly, we find that most recent models are already partially resilient to perturbations of non-causal agents but still struggle to capture indirect causal effects that involve mediator agents.
@@
To further enhance the causal robustness of the learned representations, we propose a regularization approach that seeks to preserve the causal effect of each individual neighbor in an embedding space. Specifically, we devise two variants that exploit annotations with different levels of granularity: (i) a contrastive-based regularizer using binary annotations of causal/non-causal agents; (ii) a ranking-based regularizer using continuous annotations of causal effects. Through controlled experiments, we show that both regularizers can enhance causal awareness to notable degrees. More crucially, we find that finer-grained annotations are particularly important for generalization out of the training distribution, such as higher agent density or unseen context.
@@
Finally, we introduce a sim-to-real causal transfer framework, aiming at extending the strengths of causal regularization from simulation to real-world contexts. We achieve this through cross-domain multi-task learning, i.e., jointly train the representation on the causal task in simulation and the forecasting task in the real world. Through experiments on the ETH-UCY dataset [36, 52] paired with an ORCA simulator [65], we find that the causal transfer framework enables stronger generalization in challenging settings such as low-data regimes. Despite a large gap between simulator and real-world data, our approach outperforms the AutoBots baseline with only 25% of the real data when combined with our causal regularizers. Furthermore, with just 50% of the real data, it surpasses all other models in prediction accuracy. This is achieved even in the absence of real-world causal annotations, underscoring the robustness of our framework in limited-data scenarios. We also test the causal transfer approach on the NBA dataset [70, 71], demonstrating its effectiveness in different real-world settings. As one of the first steps towards causal modeling in multi-agent systems, we hope our work sheds new light on the challenges and opportunities of learning causally-aware representations in practice.
@@ 2 Related Work
Multi-Agent Interactions. The study of multi-agent interactions has a long history. Early efforts were focused on hand-crafted rules, such as social forces [25, 48] and reciprocal collision avoidance [65, 1]. Despite remarkable results in sparse scenarios [44, 72, 21], these models often lack social awareness in more densely populated and complex environments [57]. As an alternative, recent years have witnessed a paradigm shift toward learning-based approaches, particularly the use of carefully designed neural networks to learn representations of multi-agent interactions [32]. Examples include pooling operators [2, 24, 14], attention mechanisms [68, 60, 30, 18], spatio-temporal graphs [31, 61, 37], among others [55, 8, 13]. Nevertheless, the robustness of these models remains a grand challenge [58, 56, 6]. Our work presents a solution to enhance robustness by effectively exploiting causal annotations of varying granularity.
@@
Robust Representations. The robustness of machine learning models, especially under distribution shifts, has been a long-standing concern for safety-critical applications [54]. Existing efforts have explored two main avenues to address this challenge. One line of work seeks to identify features that are invariant across distributions. Unfortunately, this approach often relies on strong assumptions about the underlying shifts [26, 39] or on access to multiple training domains [3, 34], which may not be practical in real-world settings. Another approach aims to develop models that can efficiently adapt to new distributions by updating only a small number of weight parameters, such as sub-modules [33], certain layers [35], or a small subset of neurons [10]. More recently, there has been a growing interest in exploring the potential of causal learning to address the robustness challenges [67, 15, 50, 16]. To the best of our knowledge, our work makes the first attempt in the multi-agent context, showcasing the benefits of incorporating causal relationships for stronger out-of-distribution generalization in crowded scenarios.
@@
Causal Learning. Empowering machine learning with causal reasoning has gained a growing interest in recent years [62, 63]. One line of work seeks to discover high-level causal variables from low-level observations, e.g., disentangled [12, 27, 42] or structured latent representations [43, 64]. Unfortunately, existing methods remain largely limited to simple and static datasets [41]. Another thread of work attempts to draw causal insights into dynamic decision-making [29]. In particular, recent works have proposed a couple of methods to incorporate causal invariance and structure into the design and training of forecasting models in the multi-agent context [9, 45, 40, 28, 7, 4]. However, these efforts have mainly focused on using causal implications to enhance robustness rather than explicitly examining causal relations between interactive agents. Closely related to ours, another recent study introduces a motion forecasting benchmark with annotations of causal relationships [56]. Our work addresses this benchmark’s limitations and introduces a sim-to-real approach to enhance the causal awareness of the learned representations.
3 Preliminaries
@@
In this section, we seek to formalize the robustness challenge in multi-agent representation learning through the lens of social causality. We will first revisit the design of the CausalAgents [56] benchmark, drawing attention to its caveats in labeling and evaluation. Subsequently, we will introduce a diagnostic dataset, aiming to facilitate more rigorous development and evaluation of causally-aware representations through counterfactual simulations.
@@ 3.1 Social Interaction and Social Causality
Figure 2: Illustration of indirect causal effects. The elderly pedestrian indirectly affects the trajectory of the ego agent by influencing the path of the green pedestrian. In the elderly pedestrian’s absence, the green pedestrian would proceed straight, allowing the ego agent to continue at its current speed without needing to slow down.
Figure 3: Joint effect of removing all non-causal agents in the scene, simulated in ORCA. The ego agent often changes its behavior when more than 2 non-causal neighbors are removed.
Social Interaction. Consider a motion forecasting problem where an ego agent is surrounded by a set of neighboring agents in a scene. Let $s_{t}^{i}=(x_{t}^{i},y_{t}^{i})$ denote the state of agent $i$ at time $t$ and $s_{t}=\{s_{t}^{0},s_{t}^{1},\cdots,s_{t}^{K}\}$ denote the joint state of all the agents in the scene. Without loss of generality, we index the ego agent as 00 and the rest of neighboring agents as $\mathcal{A}=\{1,2,\cdots,K\}$. Given a sequence of history observations $\bm{x}=(s_{1},\cdots,s_{t})$, the task is to predict the future trajectory of the ego agent $\bm{y}=(s_{t+1}^{0},\cdots,s_{T}^{0})$ until time $T$. Modern forecasting models are largely composed of encoder-decoder neural networks, where the encoder $f(\cdot)$ first extracts a compact representation $\bm{z}$ of the input with respect to the ego agent and the decoder $g(\cdot)$ subsequently rolls out a sequence of its future trajectory $\bm{\hat{y}}$:
@@
Social Causality. Despite remarkable progress on accuracy measures [2, 61], recent neural representations of social interactions still suffer from a significant robustness concern. For instance, recent works have shown that trajectories predicted by existing models often output colliding trajectories [39], vulnerable to adversarial perturbations [58, 5] and deteriorate under distribution shifts of spurious feature such as agent density [40]. One particular notion of robustness that has recently gained much attention is related to the causal relationships between interactive agents [56]. Ideally, neural representations of social interactions should capture the influence of neighboring agents, namely how the future trajectory of the ego agent will vary between a pair of scenes: an original scene $\bm{x}_{\varnothing}$ and a perturbed scene $\bm{x}_{R}$ where some neighboring agents $\mathcal{R}$ are removed, i.e., only the ego agent and $\mathcal{A}\setminus\mathcal{R}$ remain. We define the causal effect of the removed agents $\mathcal{R}$ as,
@@
where $\bm{y}_{\varnothing}\equiv\bm{y}$ is the future trajectory in the original scene, $\bm{y}_{\mathcal{R}}$ is the future trajectory in the perturbed scene after removing agents in $\mathcal{R}$ from scene, and $\|{\cdot}\|_{2}$ is the average point-wise Euclidean distance between two trajectories.
3.2 Caveats of CausalAgents Benchmark
While the mathematical definition of the causal effect is straightforward in the multi-agent context, measuring Eq. 2 in practice can be highly difficult. In general, it is impossible to observe a subset of agents replaying their behaviors in the same environment twice in the real world – an issue known as the impossibility of counterfactuals [53]. To mitigate the data collection difficulty, a recent benchmark CausalAgents [56] proposes another simplified labeling strategy: instead of collecting paired counterfactual scenes, it queries human labelers to divide neighboring agents into two categories: causal agents that directly influence the driving behavior of the ego agent from camera viewpoints; and non-causal agents that do not. This labeling approach is accompanied by an evaluation protocol through agent removal, assuming that robust forecasting models should be insensitive to scene-level perturbations that remove non-causal agents. Formally, the CausalAgents benchmark evaluates the robustness of a learned representation through the following measure,
@@ 3.2 Caveats of CausalAgents Benchmark
where $\mathcal{R}$ includes all or some non-causal agents in the scene.
Caveats. In spite of the precious efforts in gathering causal annotations, the CausalAgents benchmark [56] is subject to two limitations that require attention:
  1. nosep
    Annotation: The human labeling process overlooks indirect causal effects. As illustrated in Fig. 3, a neighboring agent may have a considerable causal influence on the behavior of the ego agent via mediators. The complexity of these causal chains poses a significant challenge for human annotations. Yet, these indirect effects are common in scenarios with high agent density, presenting a significant modeling challenge, which we will demonstrate in Sec. 5.
  2. nosep
    Evaluation: The evaluation protocol of the CausalAgents benchmark may inaccurately portray robustness due to oversimplification in handling agent interactions, as illustrated in  Fig. 3. This figure shows that the impact of non-causal agents grows when they are removed collectively rather than individually. While agents are marked individually as non-causal, the robustness test removes them in groups, which can exaggerate the model’s perceived robustness. Simultaneous removal of multiple non-causal agents might expose causal effects that are not detected when these agents are considered separately. For a detailed scenario illustrating this effect, please refer to Appendix C.